OpenStudy (anonymous):

Find the equation of the tangent to the curve. y=xsinx when x= pi I am so far at y=-cospi(x) +b But I need to get b, and I'm stuck. Isn't the y value just xsinx and the x value just pi?

sam (.sam.):

y'=xcosx+sinx

sam (.sam.):

dy/dx=-pi

OpenStudy (anonymous):

Yes, sorry. I mean -pi

sam (.sam.):

when x=pi, y=0

OpenStudy (anonymous):

Ah, Ok.

sam (.sam.):

\[y-0=-\pi(x-\pi)\] \[y=-\pi(x)+\pi^2\]

OpenStudy (anonymous):

Yes, thanks very much :)

sam (.sam.):

yw

OpenStudy (campbell_st):

find the deriviative f(x)' = sinx + xcosx sub x = pi \[f(\pi) = \sin(\pi) + \pi \times \cos(\pi) \] sin( pi) = 0 cos(pi) = -1 then the gradient at x = pi is m = -pi to find the y value in the point sub x = pi into the equation y = pi x sin(pi) y = 0 use point gadient formula y - 0 = -pi( x - pi)